How many ways can a coach form a team of 6 players from 15 applicants, and assign them in 6 different positions?
Answer by stanbon(75887)

How many ways can a coach form a team of 6 players from 15 applicants, and assign them in 6 different positions?
----
# of teams of 6: 15C6
---
# of ways to arrange a team: 6!
----
# of ways to do both: 15C6 * 6! = 5005*720 = 3,603,600 ways
==================
Cheers,
Stan H.